Find the equation, in slope-intercept form, of the line that passes through the points (5, 2) and (-7, 3). Write the equation in slope-intercept form.
: Find the equation, in slope-intercept form, of the line that passes through the points (5, 2) and (-7, 3). Write the equation in slope-intercept form.

Answer by jim_thompson5910(9921)


First let's find the slope of the line through the points and


m=(y[2]-y[1])/(x[2]-x[1]) Start with the slope formula.


m=(3-2)/(-7-5) Plug in y[2]=3, y[1]=2, x[2]=-7, and x[1]=5


m=(1)/(-7-5) Subtract 2 from 3 to get 1


m=(1)/(-12) Subtract 5 from -7 to get -12


m=-1/12 Reduce


So the slope of the line that goes through the points and is m=-1/12


Now let's use the point slope formula:


y-y[1]=m(x-x[1]) Start with the point slope formula


y-2=(-1/12)(x-5) Plug in m=-1/12, x[1]=5, and y[1]=2


y-2=(-1/12)x+(-1/12)(-5) Distribute


y-2=(-1/12)x+5/12 Multiply


y=(-1/12)x+5/12+2 Add 2 to both sides.


y=(-1/12)x+29/12 Combine like terms. note: If you need help with fractions, check out this solver.


y=(-1/12)x+29/12 Simplify


So the equation that goes through the points and is y=(-1/12)x+29/12